OpenStudy (anonymous):

4. Find the maximum value of e^2xy given the constraint x^2+y^2=4 . Assume that x and y are positive.

OpenStudy (anonymous):

Can anyone help me?

OpenStudy (anonymous):

\[x = \sqrt{ 4 - y^2}\] \[z = e^{2xy} = e^{2y \sqrt{4-y^2}}\] now differentiate w/ respect to y and set = 0

OpenStudy (anonymous):

I have Fx=e^2xy-2x(lamda-don't know how to type that here.) Fy=e^2xy-2y(lambda) Cannot figure out F(lamda)...

OpenStudy (anonymous):

\[ dz/dy =(2\sqrt{4-y^2} -2y\frac{y}{\sqrt{4-y^2}})e^{2y \sqrt{4 - y^2}}=0\]

OpenStudy (anonymous):

f(x,y)=e^2xy g(x,y)=x^2+y^2-4 F(x,y,lamda)=f(x,y)-lamda*g(x,y) Does any of this make sense? Fx=e^2xy-2x(lamda-don't know how to type that here.) Fy=e^2xy-2y(lambda) Cannot figure out F(lamda)...

OpenStudy (anonymous):

\[2\sqrt{4-y^2} -2y\frac{y}{\sqrt{4-y^2}} =0\] \[4 - y^2 = y^2\] \[y = \sqrt2\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

im not 100% its an educated guess

OpenStudy (anonymous):

No. but thank you! :)

OpenStudy (anonymous):

if y^2 + x^2 = 4 then x = sqrt(4 - y^2) substitute this into f(x,y) to get an explicit function f(y) then differentiate

OpenStudy (anonymous):

*a one variable function